OpenStudy (anonymous):

If it takes 526 J of energy to warm 7.40g of water by 17 degrees celsius, how much energy would be needed to warm 7.40g of water by 55 degrees celsius?

OpenStudy (anonymous):

\[Q=mc DeltaT\] Q=526 m=7.4 T=17 C= 526/17*7.4 For the answer again Q=mc DeltaT Q=7.4*55*526/17*7.4=55*526/17